OpenStudy (anonymous):

36-[1+9y-4(y+2)]=-4(3y-5)-[8(y-1)-10y+10]

OpenStudy (anonymous):

ohk so first you solve every thing in the perentheses, 36-{1+9y-4y-8] = -12y +20 - [8y -8 - 10y +10}

OpenStudy (anonymous):

36-[1+9y-4(y+2)] = -4(3y-5)-[8(y-1)-10y+10] 36- [1+9y-4y-8] = -12y+20-[8y-8-10y+10] 36-[5y-7] = -12y+20-[-2y+2] 36-5y+7 = -12y+20+2y-2 43-5y = -10y+18 5y = -25 y = -5